Setting locant numbers by locant.sty (v4.01) of XΥMTEX

Shinsaku Fujita Department of Chemistry and Materials Technology, Kyoto Institute of Technology, Matsugasaki, Sakyoku, Kyoto, 606-8585 Japan

2004/08/30

1 Introduction

1.1 Options for docstrip

option function
locant locant.sty
driver driver for this dtx file

1.2 Version Information

1 ⟨∗locant⟩

2 \typeout{XyMTeX for Drawing Chemical Structural Formulas. Version 4.01}
3 \typeout{       -- Released August 30, 2004 by Shinsaku Fujita}
4 % %%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
5 \def\j@urnalname{locant}
6 \def\versi@ndate{August 30, 2004}
7 \def\versi@nno{ver4.01}
8 \def\copyrighth@lder{SF} % Shinsaku Fujita
9 % %%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
10 \typeout{XyMTeX Macro File `\j@urnalname' (\versi@nno) <\versi@ndate>%
11 \space[\copyrighth@lder]}

2 Macros for designating locant numbers

2.1 Font Size of Locant Numbers

To change the font size of locant numbers, we can use \locantnumsize, into which a size command (\scriptsize etc.) is introduced by using the \let command. The default setting is \scriptsize as follows:

\locantnumsize
12 \let\locantnumsize=\scriptsize

2.2 Bond indicators for vertical formulas

\dbloocant \dblocant The macro \dbloocant is used to designate six edges (inner double bonds) of a six-membered ring. Any character can be asigned to the respective bond. The macro \dblocant provides a fixed set of aliphabetical charactors on the six edges.

13 \def\bdloocant#1#2#3#4#5#6{%
14 %\begin{sfpicture}(800,880)(-\shiftii,-\shifti)
15 \begin{sfpicture}(0,0)(-\shiftii,-\shifti)
16 \putratom{90}{380}{\locantnumsize #1}%
17 \putratom{200}{180}{\locantnumsize #2}%
18 \putratom{90}{-20}{\locantnumsize #3}%
19 \putlatom{-90}{380}{\locantnumsize #6}%
20 \putlatom{-200}{180}{\locantnumsize #5}%
21 \putlatom{-90}{-20}{\locantnumsize #4}%
22 \end{sfpicture}}%
23 \def\bdlocant{\bdloocant{a}{b}{c}{d}{e}{f}}

2.3 Atom indicators for vertical formulas

\sxloocant \sxlocant The macro \sxloocant is used to designate six vertices of a six-membered ring. Any character can be asigned to the respective bond. The macro \sxlocant provides a fixed set of arabic numerals on the six edges.

24 \def\sxloocant#1#2#3#4#5#6{%
25 %\begin{sfpicture}(800,880)(-\shiftii,-\shifti)%
26 \begin{sfpicture}(0,0)(-\shiftii,-\shifti)%
27 \putratom{-10}{340}{\locantnumsize #1}%
28 \putlatom{150}{260}{\locantnumsize #2}%
29 \putlatom{150}{110}{\locantnumsize #3}%
30 \putratom{-10}{40}{\locantnumsize #4}%
31 \putratom{-150}{110}{\locantnumsize #5}%
32 \putratom{-150}{260}{\locantnumsize #6}%
33 \end{sfpicture}}%
34 \def\sxlocant{\sxloocant{1}{2}{3}{4}{5}{6}}

2.4 Bond indicators for horizontal formulas

\dbloocanth \dblocanth The macro \dbloocanth is used to designate six edges (inner double bonds) of a horizontal six-membered ring. Any character can be asigned to the respective bond. The macro \dblocanth provides a fixed set of aliphabetical charactors on the six edges.

35 \def\bdloocnth#1#2#3#4#5#6{%
36 %\begin{sfpicture}(880,800)(-\shifti,-\shiftii)%
37 \begin{sfpicture}(0,0)(-\shifti,-\shiftii)%
38 \putlatom{10}{80}{\locantnumsize #1}%
39 \putratom{180}{210}{\locantnumsize #2}%
40 \putratom{380}{80}{\locantnumsize #3}%
41 \putratom{380}{-120}{\locantnumsize #4}%
42 \putratom{180}{-260}{\locantnumsize #5}%
43 \putlatom{10}{-120}{\locantnumsize #6}%
44 \end{sfpicture}}%
45 \def\bdlocnth{\bdloocnth{a}{b}{c}{d}{e}{f}}

2.5 Atom indicators for horizontal formulas

\sxloocanth \sixsghlocant The macro \sxloocanth is used to designate six vertices of a horizontal sixmembered ring. Any character can be asigned to the respective bond. The macro \sixsghlocant provides a fixed set of arabic numerals on the six edges.

46 \def\sxloocnth#1#2#3#4#5#6{%
47 %\begin{sfpicture}(880,800)(-\shifti,-\shiftii)%
48 \begin{sfpicture}(0,0)(-\shifti,-\shiftii)%
49 \putratom{40}{-20}{\locantnumsize #1}%
50 \putratom{110}{100}{\locantnumsize #2}%
51 \putlatom{290}{100}{\locantnumsize #3}%
52 \putlatom{380}{-20}{\locantnumsize #4}%
53 \putratom{110}{-140}{\locantnumsize #6}%
54 \putlatom{290}{-140}{\locantnumsize #5}%
55 \end{sfpicture}}%
56 \def\sxlocnth{\sxloocnth{1}{2}{3}{4}{5}{6}}

2.6 Locant Numbers for Cyclic Sugars

\sixsugarhloocnt \sixsugarhlocant The macro \sixsugarhloocnt is used to designate six vertices of a horizontal six-membered sugar ring. Any character can be asigned to the respective bond. The macro \sixsugarhlocant provides a fixed set of arabic numerals on the six edges.

57 \def\sixsugarhloocnt#1#2#3#4#5#6{%
58 \begin{sfpicture}(0,0)(-\shifti,-\shiftii)%
59 \putratom{40}{-20}{\locantnumsize #1}%
60 \putratom{120}{125}{\locantnumsize #2}%
61 \putlatom{400}{125}{\locantnumsize #3}%
62 \putlatom{492}{-20}{\locantnumsize #4}%
63 \putratom{120}{-160}{\locantnumsize #6}%
64 \putlatom{400}{-160}{\locantnumsize #5}%
65 \end{sfpicture}}%
66 \def\sixsugarhlocant{\sixsugarhloocnt{4}{5}{6}{1}{2}{3}}

67 ⟨/locant⟩